Suppose a worker with an annual discount rate of 0.05 percent currently resides in New South Wales and is deciding whether to remain there or to move to Victoria. There are three work periods left in the life cycle. If the worker remains in New South Wales, he will earn $20,000 per year in each of the three periods. If the worker moves to Victoria, he will earn $30,000 in each of the three periods. The cost of moving is $1000 and it would cost the worker $500 more per year to live in Victoria. If the worker decides to move she will be separated from her friends and family. The psychic costs of the workers move is estimated at $10 000. Will the worker make the move? Why or Why not?
1
Expert's answer
2019-04-17T09:38:31-0400

If the worker remains in New South Wales, then his net present value is:

NPV = - 1,000 - 10,000 + (30,000 - 500)/1.05 + (30,000 - 500)/1.05^2 + (30,000 - 500)/1.05^3 = 54,465.

If the worker moves to Victoria, then:

NPV = 30,000/1.05 + 30,000/1.05^2 + 30,000/1.05^3 = 69,336.

So, the worker will move to Victoria.